Income Statement
Year Ended December 31 (In millions) 2005 2004 2003
Net sales
Products $ 31,518 $ 30,202 $ 27,290
Service 5,695 5,324 4,534
37,213 35,526 31,824
Cost of sales
Products 27,932 27,637 25,306
Service 5,073 4,765 4,099
Unallocated coporate costs 803 914 443
33,808 33,316 29,848
3,405 2,210 1,976
Other income (expenses), net (449) (121) 43
Operating profit 2,956 2,089 2,019
Interest expense 370 425 487
Earnings before taxes 2,586 1,664 1,532
Income tax expense 761 368 479
Net earnings $ 1,825 $ 1,296 $ 1,053
Balance Sheet
December 31 (In millions) 2005 2004
Assets
Cash and cash equivalents $ 2,124 $ 1,080
Short-term investments 429 396
Receivables 4,579 4,094
Inventories 1,921 1,864
Deferred income taxes 861 982
Other current assets 495 557
Total current assets 10,409 8,973
Property, plant and equipment, net 3,924 3,599
Investments in equity securities 196 812
Goodwill 8,447 7,892
Purchased intangibles, net 560 672
Prepaid pension asset 1,360 1,030
Other assets 2,728 2,596
Total assets $ 27,624 $ 25,574
Liabilities and stockholders' equity
Accounts payable $ 1,998 $ 1,726
Customer advances and amounts in excess of costs incurred 4,331 4,028
Salaries, benefits and payroll taxes 1,475 1,346
Current maturities of long-term debt 202 15
Other current liabilities 1,422 1,451
Total current liabilities 9,428 8,566
Long-term debt 4,944 5,184
Accrued pension liabilities 1,617 1,760
Other postretirement benefit liabilities 1,277 1,236
Other liabilities 2,491 1,807
Stockholders' equity
Common stock, $1 par value per share 432 438
Additional paid-in capital 1,724 2,223
Retained earnings 7,278 5,915
Accumulated other comprehensive loss (1,553) (1,532)
Other (14) (23)
Total stockholders' equity 7,867 7,021
Total liabilities and stockholders' equity $ 27,624 $ 25,574
Consolidated Statement of Cash Flows
Year Ended December 31 (In millions) 2005 2004 2003
Operating Activities
Net earnings $ 1,825 $ 1,266 $ 1,053
Adjustments to reconcile net earnings to net cash provided by operating activities
Depreciation and amortization 555 511 480
Amortization of purchased intangibles 150 145 129
Deferred federal income taxes 24 (58) 467
Changes in operating assets and liabilities:
Receivables (390) (87) (258)
Inventories (39) 519 (94)
Accounts payable 239 288 330
Customer advances and amounts in excess of costs incurred 296 (228) (285)
Other 534 568 (13)
Net cash provided by operating activities 3,194 2,924 1,809
Investing Activities
Expenditures for property, plant and equipment (865) (769) (687)
Acquisition of business/investments in affiliated companies (784) (91) (821)
Proceeds from divestiture of businesses/Investments in affiliated companies 935 279 234
Purchase of short-term investments, net (33) (156) (240)
Other 28 29 53
Net cash used for investing activities (719) (708) (1,461)
Financing Activities
repayment of long-term debt (53) (1,069) (2,202)
Issuances of long-term debt -- -- 1,000
Long-term debt repayment and issuance costs (12) (163) (175)
Issuances of common stock 406 164 44
Repurchases of common stock (1,310) (673) (482)
Common stock dividends (462) (405) (261)
Net cash used for financing activities (1,431) (2,146) (2,076)
Net increase (decrease) in cash and cash equivalents 1,044 70 (1,728)
Cash and cash equivalents at beginning of year 1,080 1,010 2,738
Cash and cash equivalents at end of year $ 2,124 $ 1,080 $ 1,010

(a) Compute Lockheed Martin's quick ratio for 2005 and 2004. (Round your answers to two decimal places.)

2005 quick ratio = Answer ? (0.71 is the wrong answer)
2004 quick ratio = Answer ? (0.60 is the wrong answer)

(b) 2005 total debt-to-equity = Answer? (0.62/0.63 are the wrong answers)

(c) 2005 cash from operations to total debt = Answer? (0.16 is the wrong answer)

(d) 2005 free operating cash flow to total debt = Answer? (0.32/ 0.33 are wrong answers)

2004 free operating cash flow to total debt = Answer?
Answer #1

Part A

Quick ratio = (cash and cash equivalents + accounts receivable + short term investments) / current liabilities

2005 = (2124+429+4579)/9428 = 0.76

2004 = (1080+396+4094)/8566 = 0.65

Part B

total debt-to-equity = total liabilities / total stockholders' equity 2005 = (9428+4944+1617+1277+2491)/7867=19757/7867= 2.51

Part C

cash from operations to total debt = Net cash provided by operating activities / total liabilities = 3194/19757 = 0.16

Part D

free operating cash flow to total debt = free operating cash flow / total liabilities

2005 = (3194-865-784)/19757 = 0.11

2004=(2924-769-91)/(25574-7021)=0.11

Dear student,

Your answer for cash from operations to total debt is correct you can check the calculation. Detail computation of other ratios is provided. If any doubt, feel free to ask